OpenStudy (anonymous):

find the exact value of : cos(2 arctan(4/3))

OpenStudy (amistre64):

2 tan^-1(4/3) gives us an angle between (-pi/2,pi/2)...

OpenStudy (amistre64):

its a 4-3-5 triangle, I notice that

OpenStudy (amistre64):

3-4-5....but same difference lol

OpenStudy (amistre64):

the cos will equal 3/5, so 2(3/6) = 6/5 that might help us

OpenStudy (amistre64):

cos(cos^-1(6/5)) sounds right so cos(t) = 6/5 .... is that findable?

OpenStudy (amistre64):

or is that our answer :)

OpenStudy (amistre64):

to recap...for my own sanity: cos(t) = x t = 2 tan^-1(4/3) t = 2 cos^-1(3/5) t =2 cos^-1(3/5) cos(2 cos^-1(3/5)) i now this is suppose to help me....

OpenStudy (amistre64):

t/2 = cos^-1(3/5) might fit someplace

OpenStudy (amistre64):

lol...i think i confused myself again

OpenStudy (anonymous):

oh i figured it out the answer is -7/25, thanks for your help